YOU DESIGNATE ONE ROW IN YOUR GARDEN TO BROCCOLI AND PEA PLANTS. EACH BROCCOLI PLANT NEEDS 12 INCHES OF SPACE AND EACH PEA PLANT NEEDS 6 INCHES OF SPACE. THE ROW IS 10 FEET LONG. IF YOU WANT A TOTAL OF 13 PLANTS, HOW MANY OF EACH PLANT CAN YOU HAVE?
Answer by Alan3354(8285)

B = # of Br plants
P = # of Peas
B+P = 13
12B + 6P = 120 (10' = 120 inches)
B=13-P
12(13-P) + 6P = 120
156-12P+6P=120
156 - 6P = 120
-6P = -36
P = 6
B = 7